OpenStudy (jenessa_2413):

Martha, Brett, and Mike are competing in a cooking competition. They all used different amounts of oil from a can containing 8 gallons of oil. •Martha used 17% of the total oil in the can. •Brett used 0.18 of the total oil in the can. •Mike used 0.7 gallon of the total oil in the can. Who used the greatest amount of oil from the can? Show your work and explain your answer in words.

OpenStudy (jenessa_2413):

@Aveline

OpenStudy (will.h):

i got this

OpenStudy (will.h):

All of the students had 8 gallons to use and so if there are percents the covert it to real numbers lets start with Martha: she used 17% which means 0.17 and so multiply 0.17 by 8 to find the amount she used. 0.17 * 8 = 1.36 gallons Brett: 0.18 * 8 = 1.44 gallons mike: 0.7 * 8 = 5.6 gallons Now you can determine the answer i guess :)